It is known from observations that, on average, 30% of incoming solar radiation is reflected back to space, and only about 50% of it reaches the Earth
Answer #1

From law of conservation of energy:

a_\lambda I_\lambda =I_\lambda -r_\lambda I_\lambda-\tau_\lambda I_\lambda

a_\lambda+r_\lambda+\tau_\lambda =1

This is shows a process of absorption, reflection, and transmission account for all incident radiation.

When a_\lambda = 0.5Iand r_\lambda = 0.3I

Thus, \tau_\lambda =1-a_\lambda-r_\lambda = 1-0.5-0.3 = 0.2 .

The relation between transmittance is related to optical depth and to absorbance as,

T=e^{-\tau}

Therefore, the optical depth is

T=e^{-\tau} =e^{-0.2}= 0.82